OpenStudy (anonymous):

Y=2x^2-6x+c;(3/2,0.5)

OpenStudy (anonymous):

\[Y=2x^2-6x+c;(3/2,0.5)\]

OpenStudy (radar):

Use the P(3/2,0.5) x=3/2, y=.5 Plug these in to your equation to determine c

OpenStudy (anonymous):

c=5